OpenStudy (anonymous):

How do you find the cofunction of sin 65?

jimthompson5910 (jim_thompson5910):

the CO function of SINE is....

jimthompson5910 (jim_thompson5910):

any ideas what to do with that hint?

OpenStudy (anonymous):

sin (90° – x) = cos x

OpenStudy (anonymous):

Ahem.

OpenStudy (anonymous):

Csc?

jimthompson5910 (jim_thompson5910):

nope, not cosecant that's the co-function to secant

OpenStudy (anonymous):

take a look at this: http://www.mathwords.com/c/cofunction_identities.htm

OpenStudy (anonymous):

I don't understand, what does co function mean exactly?

OpenStudy (anonymous):

cos 35?

OpenStudy (anonymous):

The opposite of sin is csc

jimthompson5910 (jim_thompson5910):

close, but it's cos(25) the idea of a cofunction is that if f is some function and g(x) is the cofunction of f, then f(x) = g(90-x) or g(x) = f(90-x)

jimthompson5910 (jim_thompson5910):

examples CO SINE is the CO function of SINE (combine CO and SINE to get COSINE) So that helps you remember that cosine is the co-function of sine (or vice versa)

jimthompson5910 (jim_thompson5910):

so cos(x) = sin(90-x) and sin(x) = cos(90-x)

OpenStudy (anonymous):

Like jim said, just plug 65 into the cofunction identity for sine. sin (90 degrees- 65 degrees)= cos 25 degrees

jimthompson5910 (jim_thompson5910):

other co-function pairs are secant and cosecant tangent and cotangent

OpenStudy (anonymous):

Ok, so Sec 75=Csc 15?

OpenStudy (anonymous):

thats the cofunction?

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

:) Is there away to right it in radian form? It asks for that too

OpenStudy (anonymous):

the link i posted has the Pi form!

jimthompson5910 (jim_thompson5910):

90 degrees = pi/2 radians

jimthompson5910 (jim_thompson5910):

so just replace every occurrence of 90 with pi/2

jimthompson5910 (jim_thompson5910):

ex: cos(x) = sin(90 - x) turns into cos(x) = sin(pi/2 - x)

OpenStudy (anonymous):

So it would be sec(pi/2-75)=?

OpenStudy (anonymous):

There is nothing to do within the parenthesis

jimthompson5910 (jim_thompson5910):

you need to convert 75 degrees to radians, otherwise it's not the same

OpenStudy (anonymous):

5(pi)/12

OpenStudy (anonymous):

pi/2-5(pi)/12

jimthompson5910 (jim_thompson5910):

good, now just combine those fractions

OpenStudy (anonymous):

pi/12?

jimthompson5910 (jim_thompson5910):

good, you got it

jimthompson5910 (jim_thompson5910):

so sec(75) = csc(15) becomes sec(5pi/12) = csc(pi/12)

OpenStudy (anonymous):

Just like that! I got it, thanks a lot :D

jimthompson5910 (jim_thompson5910):

yw